How many Liters of a 40% solution and a 16% solution must be mixed to obtain 20 liters of a 22% solution? from there i got the equation .4x + .16(20-x)= 20(.22)

from .4x+ .16(20-x)= 20 (.22) how is the answer x=5 and is 5 liters at 40% and 15 liters at 16%

need a break down of how to get x=5, 5 liters at 40 %, and 15 liters at 16%

Answer by rfer(10432)

.4x+.16(20-x)=20(.22)
.4x+3.2-.16x=4.4
.20x=1.2
x=6 lt of 40%
20-x=14 lt of 16%